Newham Corporation produces and sells two products. In the most recent month, Product R10L had sales of $28000 and variable expenses of $6440. Product X96N had sales of $22000 and variable expenses of $7560. The fixed expenses of the entire company were $32,710. The break even point for the entire company is closest to:

A.). $32,710

B.) $17,290

C.) $45,431

D.) $46,710
Answer #1

Total sales = 28,000 + 22,000 = 50,000

Total Variable costs = 6440 + 7560 = 14,000

Contribution margin = 50,000 - 14,000 = 36,000

CM ratio = 36,000/50,000 = 78%

Breakeven point = Fixed cost /CM ratio

= 32,710/72%

= 45,431